Property Tax Rate and Revenues

Steuben County anticipates appropriations of $10,500,000 and revenue from the state and other sources of $900,000. During the last fiscal year, the tax rate was $0.85 per $100 of assessed valuation, applied to property assessed at $1,000 million. Uncollectible property taxes are negligible. In the current year, the county’s reassessment program increased the aggregate assessed value of property to $1,050 million.

Required

Enter numerical answers using all zeros (for parts a & c). Do not abbreviate answers to millions or thousands.

a. Compute the property tax revenue raised during the last fiscal year.

$Answer

b. Compute the tax rate needed to balance the current budget.

Round answer to three decimal places.

$Answer

c. Assume that the Steuben County charter requires a balanced budget and that increasing the tax rate is politically unacceptable. Compute the maximum amount of appropriations that can be funded by property taxes given the increase in assessed valuation.

$Answer
Answer #1

PART A

Property Tax revenue raised during the last fiscal year is:

Amount in Dollar = 1,000 x 1,000,000 = $1,000,000,000

Property Tax = (1,000,000,000) x (0.85/100) = $ 8,500,000.

Therefore total property tax revenue raised in the last fiscal year is $8,500,000

PART C

Assuming the revenue from state and other sources amounts to 900,000 in the current year as well.

Total amount of appropriations that can be funded on increased property valuation is :

Amount in Dollar = 1,050 x 1,000,000 = $ 1,050,000,000

Property Tax = ( 1,050,000,000) x (0.85/100) = $ 8,925,000

Total Appropriation = 8,925,000 + 900,000 = $ 9,825,000